### v2

# Robust Identity Perceptual Watermark Against Deepfake Face Swapping

Abstract
Notwithstanding offering convenience and entertainment to society, Deepfake face swapping has caused critical privacy issues with the rapid development of deep generative models. Due to imperceptible artifacts in high-quality synthetic images, passive detection models against face swapping in recent years usually suffer performance damping regarding the generalizability issue. Therefore, several studies have been attempted to proactively protect the original images against malicious manipulations by inserting invisible signals in advance. However, the existing proactive defense approaches demonstrate unsatisfactory results with respect to visual quality, detection accuracy, and source tracing ability. In this study, to fulfill the research gap, we propose the first robust identity perceptual watermarking framework that concurrently performs detection and source tracing against Deepfake face swapping proactively. We assign identity semantics regarding the image contents to the watermarks and devise an unpredictable and nonreversible chaotic encryption system to ensure watermark confidentiality. The watermarks are encoded and recovered by jointly training an encoder-decoder framework along with adversarial image manipulations. Falsification and source tracing are accomplished by justifying the consistency between the content-matched identity perceptual watermark and the recovered robust watermark from the image. Extensive experiments demonstrate state-of-the-art detection performance on Deepfake face swapping under both cross-dataset and cross-manipulation settings.
1 Introduction
The prevalent Deepfake face swapping technique has received widespread attention with the development of deep generative models. By swapping the facial identity from the source image onto the target one, Deepfake face swapping can bring convenience to people’s lives and benefit society from the perspective of human entertainment such as film-making. Nevertheless, misusing the technique has brought privacy issues to various groups of victims including celebrities, politicians, and even every human being [47]. Therefore, it is necessary to prevent malicious attacks while maintaining positive usages of Deepfake face swapping.
Most existing countermoves [38, 22, 60, 46, 13, 33, 50, 35] regarding Deepfake face swapping follow the pipeline of devising a deep detection model that passively investigates and analyzes the underlying synthetic artifacts to distinguish the fake images. While accomplishing progresses in the early stages, passive detection suffers the generalization challenge such that obvious performance damping occurs when attempting to detect fake materials produced by unknown synthetic algorithms [45]. Moreover, following rapid improvements in deep generative algorithms, it has become burdensome to locate perceivable manipulation traces within high-quality synthetic images. To overcome this issue, several recent studies consider executing proactive defense against synthetic manipulations. In other words, the goal is to protect the original images by inserting imperceptible signals in advance of posting in public and being manipulated by Deepfake unexpectedly.
The proactive approaches can be categorized into two main strategies, namely, distorting [15, 44, 48] and watermarking [55, 31, 3, 30, 25]. The former learns a distortion that nullifies potential synthetic manipulations by inserting it into the original images, and the latter encodes semi-fragile watermarks into original images and determines the authenticity based on the existence of watermarks afterward. Both ways aim to proactively protect the original images without affecting the visual qualities. While demonstrating preliminary initiations, the existing approaches suffer several flaws. Firstly, in the current research stage, both the distorting and watermarking approaches lack generalization ability on unseen datasets when facing unknown synthetic manipulations. Secondly, although the semi-fragile watermarks can help determine real and fake based on their existence, they are prone to be unstable when facing unavoidable common image processing operations. Moreover, watermarks that are fragile to malicious synthetic manipulations cannot help conduct source tracing on the victim target images simultaneously. In particular, in an integral chain of evidence for the forensic investigation in the cybercrime cases related to Deepfake face swapping, besides correctly addressing the falsification, it is crucial to retrieve intelligence information of the source materials [1, 20, 51]. In other words, tracing the original target images along with the detection task can help fulfill the evidence integrality. Lastly, while preventing privacy issues, directly adding distortions not only unavoidably degrades the visual qualities of the images [11], but also unfavorably disables benign adoptions of Deepfake face swapping.
To address the aforementioned issues, in this study, we propose a novel robust identity perceptual watermarking framework that promotes proactive defense against Deepfake face swapping. In general, we establish semantically meaningful identity perceptual watermarks that are robust and invisible based on image contents and embed them into original images for detection and source tracing purposes simultaneously, while allowing benign usages of face swapping algorithms. First, based on the identities of facial images, we construct securely protected identity perceptual watermarks by introducing a chaotic encryption system. Then, we devise an end-to-end robust watermarking framework that consists of an encoder-decoder architecture to be trained with common and face swapping manipulation pools. Relying on the promising watermark recovery performance, the authenticity of a candidate image is determined by the consistency between the content-matched identity perceptual watermark and the robustly recovered watermark from the image. The source tracing is simultaneously achieved with respect to the identity semantics carried by the recovered watermark. Extensive experiments prove the generalization ability of our proposed approach under cross-dataset and cross-manipulation settings with average watermark recovery accuracies all above 96%, strongly outperforming the contrastive methods. The outstanding detection performance against Deepfake face swapping with AUC scores all above 97% also consistently outperforms the state-of-the-art passive Deepfake detectors.
Our contributions can be summarized as follows:
  • We propose a novel idea of identity perceptual watermarks based on image contents to proactively defend against Deepfake face swapping regarding content-watermark consistencies. A chaotic encryption system is collaboratively devised to ensure watermark confidentiality.
  • We present an encoder-decoder framework to robustly and invisibly embed the watermarks into facial images. To the best of our knowledge, we are the first to concurrently achieve detection and source tracing for Deepfake face swapping by embedding a single watermark.
  • Experiments qualitatively and quantitatively demonstrate the promising watermark recovery accuracy and Deepfake face swapping detection performance of our method along with outstanding generalization ability, outperforming the related state-of-the-art algorithms.
2 Related Work
Existing forensic investigations on Deepfake face swapping mostly rely on devising passive detectors [38, 22, 60, 13, 32, 33, 50] that analyze traces of suspected images in various feature domains. In view of the rising challenge prompted by the improving synthetic quality, instead of searching for synthetic artifacts, proactive defense aims to protect original images by inserting imperceptible signals before potential manipulations. In other words, while malicious face swapping happens after the target image is publicly posted (e.g., on Twitter or Instagram), proactive actions are performed in advance in two ways: distortions and watermarks.
2.1 Proactive Distortions
Proactive distortions [37, 17, 57, 56, 2] are usually inserted into pristine images to nullify the synthetic models following $\tilde{x}=x+\eta$, where $x$ and $\tilde{x}$ are the pristine images before and after the learned invisible distortion $\eta$ is added, respectively. In particular, recently, Anti-Forgery [44] exploits the vulnerability of GAN-based models and supervisely learns a particular distortion in the lab color space for a specific pristine image regarding each synthetic algorithm. CMUA [15] presents a two-level perturbation fusion strategy and heuristically learns a universal distortion that can disrupt multiple synthetic algorithms. TAFIM [2] introduces image-specific perturbations with better robustness via attention-based fusion and refinement. Besides adding invisible perturbations that could heavily distort the synthetic results, Wang et al. [48] further promoted a detector that distinguishes the distorted images with high accuracy since machines do not share similar perspectives as human eyes.
2.2 Proactive Watermarks
Digital watermarks [41] are originally used for copyright protection and authentication [54, 9, 23], and they are thus designed to have promising robustness. Ever since the occurrence of the first deep neural network (DNN) based end-to-end trainable watermarking framework, HiDDeN [61], various follow-up algorithms [18, 28, 40, 8, 16] emerge based on the encoder-decoder architecture for robust watermarking. The encoder reconstructs a visually identical image by inserting a secret message into the original one, and the decoder takes charge of recovering the encoded message from the watermarked image. Particularly, MBRS [18] proposes a novel training strategy within mini-batches to enhance the JPEG robustness of watermark embedding. To simultaneously improve watermark imperceptibility and robustness against unknown image noises, CIN [28] combines the invertible and non-invertible mechanisms in the watermark embedding and recovery pipelines. ARWGAN [16] enforces more abundant features from images for robust watermarking such that the image visual quality and watermark recovery accuracy can be concurrently boosted.
Figure 1: Framework of the proposed approach. An original clean image $I$ is first passed to generator $G_{m}$ to generate the corresponding identity perceptual watermark $m$. Then, $I$ and $m$ are fed to the encoder for watermark embedding. Manipulation pools $P_{\textrm{common}}$ and $P_{\textrm{swap}}$ help preserve watermark robustness during supervision. A discriminator helps maintain the visual quality. In the end, the watermark $m_{\textrm{rec}}$ is recovered by passing the manipulated image to the decoder.
Regarding the relatively novel term of proactive defense, few approaches have been attempted with watermarks. Yu et al. [58, 59] assigned particular fingerprints to images via robust watermarking regarding generative models for the attribution purpose. Different from the robust watermarking studies, FaceGuard [55] and FaceSigns [31] propose semi-fragile watermarks such that they stay robust to benign image post-processing operations but become fragile when facing malicious face manipulations. The authors claimed to distinguish fake materials if no watermark is detected in candidate images. Recently, a separable watermarking framework, SepMark [51], is conducted with two watermark decoders, a Tracer and a Detector, where the former recovers robust watermarks for source tracing, and the latter extracts semi-fragile watermarks to determine real and fake. To the best of our knowledge, robust watermarking has been barely attempted to detect Deepfake face swapping adaptively, and performing detection and source tracing simultaneously with a single watermark remains largely unexplored in the current research domain.
3 Methodology
3.1 Overview
In this paper, we propose an identity perceptual watermarking framework to adaptively embed imperceptible robust watermarks into images and proactively achieve detection and source tracing against Deepfake face swapping by precisely recovering the encoded watermarks. The model framework is demonstrated in Fig. 1. First, we devise a watermark generator $G_{m}$ to securely prepare the binary watermarks by assigning identity perceptual semantics with respect to the image contents. Then, the encoder is fed with the clean image $I$ and the corresponding watermark $m$ for encoding. To guarantee robustness, we apply a common manipulation pool $P_{\textrm{common}}$ and a face swapping manipulation pool $P_{\textrm{swap}}$ as adversaries to the watermarked image $I_{\textrm{rec}}$ before recovering the watermark $m_{\textrm{rec}}$ via the decoder. Meanwhile, a discriminator that distinguishes the watermarked and clean images is adopted in the training stage to enhance image visual quality. By assigning constraints as discussed in Section 3.5, the framework is trained to satisfy our objectives.
In practice, for an original image $I$ with facial identity $t$, the identity perceptual watermark $m$ is produced by $G_{m}$ and embedded into $I$ to generate the watermarked image $I_{\textrm{rec}}$ via the encoder. Upon the malicious face swapping manipulation, a synthetic facial image $I_{\textrm{fake}}$ is derived by swapping identity $s$ onto $I_{\textrm{rec}}$. By utilizing the decoder, the watermark $m_{\textrm{rec}}$ which is faithfully consistent with $m$ can be robustly recovered from $I_{\textrm{fake}}$. Since $I_{\textrm{fake}}$ contains facial identity $s$, its corresponding watermark $m_{s}$ can be generated using $G_{m}$, and demonstrating the inconsistency between $m_{\textrm{rec}}$ and $m_{s}$ can accomplish the detection task on Deepfake face swapping. On the other hand, as $m_{\textrm{rec}}$ is robustly recovered, the trusted original image $I$ with identity perceptual watermark $m$ can be traced by inspecting the matchings between identities and watermarks in an officially recorded hash map.
3.2 Identity Perceptual Watermarks
Unlike early frameworks that enforce random initialization, we derive the watermark for each facial image by analyzing the identity semantics. In particular, we propose an identity perceptual watermark generator $G_{m}$ that generates binary watermark $m$ with length $l$ based on the facial identity of image $I$. Inspired by hashing algorithms [36, 24], the pipeline of $G_{m}$ follows two critical characteristics: collision free and one-way computational.
Collision Free. Since the goal is to detect Deepfake face swapping by addressing inconsistencies between the recovered and content-matched watermarks of candidate images, it is necessary to ensure that no two identical watermarks belong to different facial identities. Face recognition, a task to identify different facial identities by producing distinguishable embeddings, has been proved the ability on various large datasets with reliable models. In this study, impressed by the outstanding achievements, we adopt well-performed face recognition algorithms that retrieve identity embeddings in the form of high-dimensional vectors. To match up the required watermark length $l$, we first select a large dataset that contains sufficient unique identities and retrieve the identity embeddings $E$ with dimension $d_{E}$ accordingly. Then, we apply the principal component analysis (PCA) [34] to $E$ and learn a transform rule $\phi$ that projects the identity embeddings of the images from dimension $d_{E}$ down to $l$ while optimally preserving the original correlations between facial identities with respect to the principal components. Thereafter, for the purpose of constructing binary watermarks, we conduct a unit-variance scaling on the projected embeddings $E_{\textrm{pca}}$ to the range between 0 and 1 by
$E_{\textrm{scale}}=\frac{E_{\textrm{pca}}-\min(E_{\textrm{pca}})}{\max(E_{\textrm{pca}})-\min(E_{\textrm{pca}})},$ (1)
where $E_{\textrm{scale}}$ denotes the scaled embeddings with respect to the standard deviation, which favorably preserves the distribution of $E_{\textrm{pca}}$. In the end, we set a cutoff value $0\leq c\leq 1$ to round the entries of $E_{\textrm{scale}}$ to binary values accordingly and obtain the identity perceptual watermarks $E_{\textrm{binary}}$. Besides, a watermark collision check is arranged in Section 4.4 to validate the collision-free characteristic.
Figure 2: Visual effects of common image manipulations on the watermarked images. The first two columns refer to the raw and watermarked images, and each of the rest rows displays the visual effects of a manipulation algorithm on the watermarked images.
One-Way Computational. Since we concurrently conduct detection and source tracing, an adversary with sufficient knowledge of our work may aim to replace the encoded watermark with a content-matched one. To prevent hostile attacks [42, 6], we enforce the watermark to be one-way computational, that is, it is insufficient to recover the watermark construction pipeline by merely acquiring the watermark from the image. Specifically, besides keeping $\phi$ encapsulated, we design a chaotic encryption [29] system on $E_{\textrm{binary}}$ using a logistic map
$x_{i+1}=rx_{i}(1-x_{i}),$ (2)
where constants $x_{0}$ at $i=0$ and $r$ are the initial condition and control parameter, respectively. $x_{i}$ denotes the $i$-th value in the chaotic map and $0<=i<l$. The values in $X=\{x_{1},x_{2},...,x_{l}\}$ are derived accordingly, and a binary encryption key $K=\{k_{1},k_{2},...,k_{l}\}$ is computed via
$k_{i}=\lfloor x_{i}p^{q}\bmod{2}\rfloor,$ (3)
where $p$ and $q$ are pre-defined prime constants. Finally, the encrypted identity perceptual watermark $m$ for image $I$ is derived by a logical exclusive OR (XOR) operation between $K$ and the unencrypted watermark from $E_{\textrm{binary}}$. Relying on the sensitive dependence on the initial condition and chaotic map nonlinearity characteristics, the final watermarks leveraged in the watermarking framework are unpredictable and nonreversible unless having access to Eqn.(2) and Eqn.(3) including the critical constant coefficients. Specifically, the entry values in $X$ vary drastically with respect to the values of $x_{0}$ and $r$, and the watermarks are thus securely protected. On the other hand, with authority on $K$, performing an extra XOR operation on $m$ using $K$ can easily decrypt the watermarks back to the ones in $E_{\textrm{scale}}$ to trace back for the identity information.
Figure 3: Visual effects of Deepfake face swapping image manipulations on the watermarked images. The first column refers to the source images that provide facial identity for face swapping. Columns 2 and 3 present the raw and watermarked target images. Every two columns of the remaining exhibit the face swapping visualizations on the raw and watermarked target images.
3.3 Watermark Encoding
For an input image $I$, we insert the watermark $m$ using an encoder that is mainly based on convolutional neural networks (CNN) with squeeze-and-excitation networks (SENet) [12] and diffusion blocks. Specifically, the image features and watermark features are separately analyzed and then concatenated to reconstruct the watermarked image $I_{\textrm{rec}}$.
Image Features. In this paper, we preserve the width and height of $I$ in the propagation process to avoid information loss in image reconstruction. We first feed $I$ to a convolutional block (denoted as Conv. Block) that contains a CNN layer, a batch normalization layer, and a ReLU activation function in a sequential order to exaggerate the number of feature channels. Squeeze-and-excitation network (SENet) [12], an on-the-shelf mechanism that improves channel interdependencies, can favorably perform feature recalibration on the input image. As depicted in Fig. 1, $N_{\textrm{enc}}$ repeated SENets (denoted as SE Block) are applied to seek proper channels and elements within the feature map for watermark encoding without jeopardizing the visual quality.
Watermark Features. As for the binary identity perceptual watermark $m$ of length $l$, we first increase the length linearly via a fully connected layer followed by dimension expansion via shape rearrangement and then leverage a convolutional block to elaborate the channel perspective. After that, we devise $N_{\textrm{enc}}$ consecutive diffusion blocks (denoted as Diff. Block) utilizing deconvolutions and smoothly diffuse the feature elements that represent $m$ to the same dimension as the width and height of $I$. Similarly, a series of $N_{\textrm{enc}}$ SE Blocks is promoted for the preparation of watermarking.
Image Reconstruction. Lastly, features of $I$ and $m$ are concatenated with respect to the channel dimension followed by a convolutional block and a vanilla CNN layer to reconstruct the image $I_{\textrm{rec}}$ that is visually identical to $I$.
3.4 Watermark Recovery
Manipulation Pools. In real-life scenarios, images inevitably suffer degradation by common post-processing operations and noises such as compression and blurring upon uploading and spreading on social media. Therefore, to satisfy the objectives, as shown in Fig. 1, we construct two manipulation pools, $P_{\textrm{common}}$ and $P_{\textrm{swap}}$, to supervise the proposed framework. On the one hand, the common manipulation pool $P_{\textrm{common}}$ includes the benign post-processing operations that are commonly seen in the real world. On the other hand, we enclose Deepfake face swapping models in the face swapping pool $P_{\textrm{swap}}$ to preserve watermark robustness when facing malicious attacks.
Decoder. We feed the post-manipulation image to the decoder for watermark recovery. The decoder follows a reverse pipeline of the encoder. In particular, after extracting features back to a large number of channels, we apply $N_{\textrm{se}}$ SE Blocks, where $N_{\textrm{se}}=N_{\textrm{dec}}+3$, to gradually expand the number of channels and seek the determinant elements that hide the watermark information. Then, $N_{\textrm{dec}}$ SE Blocks without channel modification are performed after narrowing down the number of channels. In the end, a convolutional block, a flattening operation, and a linear projection are sequentially applied to recover the watermark $m_{\textrm{rec}}$.
3.5 Objective Functions
During the training stage, we employ a total of 4 loss functions for guidance, namely, reconstruction loss, recovery loss, adversarial loss, and generative loss.
Reconstruction Loss. We embed the watermarks into images and pursue the uttermost similarity between the original and watermarked images by assigning a pixel-wise $L_{2}$ constraint on the encoder,
$L_{\textrm{enc}}=\lVert\textrm{Enc}(\theta_{\textrm{enc}},I,m)-I\rVert_{2},$ (4)
where $I$ denotes the clean image before watermarking, $m$ represents the watermark, and $\theta_{\textrm{enc}}$ encloses the parameters of the encoder $\textrm{Enc}(\cdot)$.
Recovery Loss. Similarly, a bit-wise $L_{2}$ constraint is applied on the decoder for watermark recovery,
$L_{\textrm{dec}}=\lVert\textrm{Dec}(\theta_{\textrm{dec}},\textrm{Enc}(\theta_{\textrm{enc}},I,m))-m\rVert_{2},$ (5)
where $\theta_{\textrm{dec}}$ refers to the parameters of the decoder $\textrm{Dec}(\cdot)$. At the same time, the encoder is updated alongside.
Adversarial Loss. Following the convention of robust watermarking [61, 18, 28], a discriminator $D$ with parameters $\theta_{d}$ is applied to improve visual qualities of the images following
$L_{\textrm{adv}}=-\mathbb{E}(\log(D(\theta_{d},I)))+\mathbb{E}(\log(1-D(\theta_{d},\textrm{Enc}(\theta_{\textrm{enc}},I,m)))).$ (6)
The discriminator is implemented as a binary classifier with $N_{\textrm{dis}}$ convolutional blocks, and it adversarially learns to identify clean and watermarked images on each pair of $I$ and $I_{\textrm{rec}}$ during model supervision.
Table 1: Quantitative visual quality evaluation of the watermarked images on CelebA-HQ. Information includes model name, image resolution, watermark length, PSNR, and SSIM. The best performance for each resolution is in bold and the second best performance is marked with underscore.
[…]
    Statistics adopted from the published paper due to unavailable source code for reproduction.
Generative Loss. When training with the face swapping manipulation pool $P_{\textrm{swap}}$, we arrange an extra objective using an $L_{2}$ constraint to make sure the embedded watermarks do not affect the synthetic results following
$L_{\textrm{gen}}=\lVert G(I,I_{s})-G(\textrm{Enc}(\theta_{\textrm{enc}},I,m),I_{s})\rVert_{2},$ (7)
where $I_{s}$ is the image that provides identity information for the face swapping model $G(\cdot)$.
Overall Objective Function. Coefficients $\lambda_{\textrm{enc}}$, $\lambda_{\textrm{dec}}$, $\lambda_{\textrm{adv}}$, and $\lambda_{\textrm{gen}}$ are assigned to compute the weighted sum in the overall objectives $L_{\textrm{common}}$ and $L_{\textrm{swap}}$ for $P_{\textrm{common}}$ and $P_{\textrm{swap}}$ following
$L_{\textrm{common}}=\lambda_{\textrm{enc}}L_{\textrm{enc}}+\lambda_{\textrm{dec}}L_{\textrm{dec}}+\lambda_{\textrm{adv}}L_{\textrm{adv}},$ (8)
and
$L_{\textrm{swap}}=L_{\textrm{common}}+\lambda_{\textrm{gen}}L_{\textrm{gen}}.$ (9)
4 Experiment
4.1 Implementation Details
Datasets. Experiments are conducted on the high-quality CelebA-HQ dataset [19] with 30,000 image samples and 6,217 unique identities following the official split [26] for training, validation, and testing. To further validate the generalization ability, we tested the trained model on the Labeled Faces in the Wild (LFW) dataset [14] with 5,749 unique identities.
Manipulation Pools. $P_{\textrm{common}}$ consists of the following manipulations for both training and testing: Dropout, Resize, Jpeg Compression, Gaussian Noise, Salt and Pepper, Gaussian Blur, and Median Blur. $P_{\textrm{swap}}$ includes SimSwap [5] solely in the training phase and adopts SimSwap, InfoSwap [10], UniFace [52], and E4S [27] with state-of-the-art performance for cross-manipulation evaluation in the testing phase. Since there are various common manipulations to be considered and they only bring noises without modifying the facial contents, we first pre-trained the model using $P_{\textrm{common}}$ for sufficient iterations. Thenceforth, we slowly fine-tuned it with the help of $P_{\textrm{swap}}$ by introducing random images that provide different facial identities for swapping onto $I_{\textrm{rec}}$.
Parameters. Upon watermark generation, CelebA-HQ is adopted to derive the transform rule $\phi$ with cutoff value $c=0.5$, and we set $x_{0}=0.1$, $r=3.93$, $p=5$, and $q=11$ for the constant coefficients in Eqn.(2) and Eqn.(3). We conducted watermark encoding with lengths $64$ and $128$ for images at $128\times 128$ and $256\times 256$ resolutions, respectively, and we set $N_{\textrm{enc}}$ to $3$ and $4$, $N_{\textrm{dec}}$ to $1$ and $2$, and $N_{\textrm{dis}}$ to $3$ and $4$ for the two resolutions. In this study, we trained the model in two steps. The proposed approach is firstly trained with $P_{\textrm{common}}$ by setting $\lambda_{\textrm{enc}}=1$, $\lambda_{\textrm{dec}}=10$, and $\lambda_{\textrm{adv}}=0.01$ with a learning rate of $31e-$ 3 to guarantee the promising watermark recovery performance regarding common manipulations. Then, the trained model is further tuned on SimSwap by setting $\lambda_{\textrm{enc}}=10$, $\lambda_{\textrm{dec}}=1$, $\lambda_{\textrm{adv}}=0.01$, and $\lambda_{\textrm{gen}}=5$ with a lower learning rate of $51e-$ 5. This process mildly adjusts the framework to stay robust when facing both common and face swapping manipulations. Adam optimizers [21] are adopted through the entire training pipeline for each model component. Our model is implemented using PyTorch on $4$ Tesla V100 GPUs with a batch size of $32$.
Table 2: Quantitative comparison on CelebA-HQ at 128 and 256 resolutions regarding the bit-wise recovery accuracy of the watermarks under common manipulations and malicious face swapping manipulations. The best performance for each resolution is in bold and the second best performance is marked with underscore.
Manipulations HiDDeN  [61] $128\times 128$ MBRS  [18] $128\times 128$ RDA  [58] $128\times 128$ CIN  [28] $128\times 128$ ARWGAN  [16] $128\times 128$ Ours $128\times 128$ MBRS  [18] $128\times 128$ FaceSigns  [31] $128\times 128$ Ours $256\times 256$
Dropout 82.44% 100.00% 94.76% 100.00% 97.25% 99.99% 99.19% 55.36% 99.99%
Resize 82.01% 100.00% 99.94% 99.17% 93.73% 100.00% 100.00% 98.63% 99.08%
Jpeg 67.84% 99.49% 66.85% 96.86% 57.42% 99.48% 99.69% 82.64% 99.99%
GaussianNoise 51.36% 99.60% 60.18% 86.00% 53.45% 98.63% 58.31% 53.61% 99.97%
SaltPepper 52.30% 99.37% 66.62% 93.95% 64.55% 95.54% 65.33% 73.95% 97.54%
GaussianBlur 73.04% 100.00% 99.95% 100.00% 85.22% 100.00% 72.05% 98.68% 97.93%
MedBlur 82.72% 100.00% 99.98% 97.03% 96.66% 100.00% 98.06% 99.86% 99.01%
Average 70.24% 99.78% 84.04% 96.14% 78.33% 99.09% 84.66% 80.39% 99.07%
SimSwap [5] 50.02% 49.98% 50.00% 50.28% 52.06% 99.99% 50.00% 49.74% 99.98%
InfoSwap [10] 50.07% 50.82% 50.01% 50.60% 47.94% 99.31% 50.71% 50.00% 99.73%
UniFace [52] 54.98% 50.22% 71.15% 46.01% 59.30% 99.57% 49.98% 50.59% 94.86%
E4S [27] 49.19% 50.07% 63.03% 50.55% 49.81% 93.19% 50.07% 49.73% 93.33%
Average 51.07% 50.27% 58.55% 49.36% 52.28% 98.02% 50.19% 50.02% 96.98%
Table 3: Quantitative visual quality evaluation of the watermarked images on LFW. Information includes model name, image resolution, watermark length, PSNR, and SSIM. The best performance for each resolution is in bold and the second best performance is marked with underscore.
[…]
HiDDeN [61] $128\times 128$ 30 33.27 0.883
MBRS [18] $128\times 128$ 30 32.78 0.761
RDA [58] $128\times 128$ 100 42.71 0.959
CIN [28] $128\times 128$ 30 42.89 0.982
ARWGAN [16] $128\times 128$ 30 39.94 0.926
Ours $128\times 128$ 64 46.00 0.993
MBRS [18] $256\times 256$ 256 45.37 0.969
FaceSigns [31] $256\times 256$ 128 38.23 0.910
Ours $256\times 256$ 128 45.42 0.994
4.2 Experiment on CelebA-HQ
In this section, experiments are conducted on the testing set of CelebA-HQ. Since no existing work has performed detection and source tracing on face swapping images using a single watermark to our knowledge, we selected the robust watermarking frameworks (HiDDeN [61], MBRS [18], and CIN [28]), proactive robust Deepfake watermarking frameworks (RDA [58], SepMark [51], and ARWGAN [16]), proactive semi-fragile Deepfake watermarking frameworks (FaceSigns [31]), and proactive Deepfake distortion frameworks (AntiForgery [44] and CMUA [15]) for comparison. For fairness, the publicly available trained model weights that correspond to optimal performance are directly adopted. Statistics of the recent proactive separable watermark framework, SepMark [^1] SepMark [51] is only compared in Table 1 because the reported results are on CelebA-HQ, but the parameters are unrevealed for the common manipulations, leading to unavailability in maintaining a fair comparison in the remaining experiments. [51], are copied from the published paper due to unavailable source code.
4.2.1 Qualitative Evaluation
We visualized randomly drafted images to validate the visual quality of the watermarked images. As exhibited in Fig. 2, despite knowing the existence of the watermarks, differences between the raw and watermarked images in the first and second columns, respectively, are visually imperceptible. Starting from column 3, we illustrated the visual effects of the common manipulations on the watermarked images, one for each column. In Fig. 3, the face swapping effects are exhibited in every two columns starting from column 4. The source images that provide the desired synthetic facial identities for each row are placed in the first column. Columns 2 and 3 display the raw and watermarked target images, while every pair of the remaining columns illustrates the face swapping results on the raw and watermarked target images, respectively. It is worth noting that no distortion can be observed in the face swapping results using watermarked target images. In other words, the encoded watermarks do not affect the visual qualities of the synthetic faces, preserving the possible benign usages of Deepfake face swapping in the industry.
4.2.2 Quantitative Evaluation
In this section, we quantitatively evaluated the visual qualities of watermarked images and the watermark recovery accuracies after common and face swapping manipulations.
Visual Quality. We employed the average peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) [49] between the raw and watermarked images to evaluate the visual quality of watermarked images. As Table 1 reports, while the goal of embedding watermarks into raw images imperceivably stands for all contrastive methods, our proposed approach achieves the state-of-the-art visual quality in terms of PSNR and SSIM at both resolution levels. Meanwhile, it can be observed that although pursuing the same objective in visual quality, the proactive distortion models, AntiForgery [44] and CMUA [15], have revealed relatively unsatisfactory performance since the noises and perturbations are directly added to raw images while watermarking approaches execute image reconstructions.
Table 4: Quantitative comparison on LFW at 128 and 256 resolutions regarding the bit-wise recovery accuracy of the watermarks under common manipulations and malicious face swapping manipulations. The best performance for each resolution is in bold and the second best performance is marked with underscore.
Manipulations HiDDeN  [61] $128\times 128$ MBRS  [18] $128\times 128$ RDA  [58] $128\times 128$ CIN  [28] $128\times 128$ ARWGAN  [16] $128\times 128$ Ours $128\times 128$ MBRS  [18] $128\times 128$ FaceSigns  [31] $128\times 128$ Ours $256\times 256$
Dropout 82.84% 99.99% 89.72% 100.00% 97.28% 99.96% 99.30% 56.45% 99.97%
Resize 81.97% 100.00% 99.89% 99.98% 95.07% 99.99% 100.00% 98.91% 98.16%
Jpeg 56.17% 99.42% 66.77% 97.06% 70.86% 99.38% 99.08% 75.64% 99.99%
GaussianNoise 51.30% 99.80% 60.83% 87.54% 53.11% 99.63% 55.61% 59.90% 99.92%
SaltPepper 52.93% 99.63% 67.31% 91.71% 62.96% 95.89% 61.15% 67.99% 95.99%
GaussianBlur 72.89% 100.00% 99.88% 99.99% 86.39% 99.99% 73.93% 99.01% 97.11%
MedBlur 82.35% 99.99% 99.98% 99.98% 96.93% 99.99% 99.38% 99.91% 98.13%
Average 68.63% 99.83% 83.48% 96.61% 80.37% 99.26% 84.06% 79.69% 98.47%
SimSwap [5] 49.96% 49.92% 50.04% 50.25% 50.00% 99.37% 50.04% 49.77% 99.96%
InfoSwap [10] 50.06% 50.04% 50.10% 50.95% 50.86% 98.69% 50.23% 49.91% 98.67%
UniFace [52] 53.74% 49.78% 70.37% 49.92% 59.16% 98.48% 50.18% 52.25% 97.06%
E4S [27] 49.34% 50.08% 67.78% 50.22% 48.64% 97.27% 49.99% 50.15% 97.00%
Average 50.77% 49.96% 59.57% 50.33% 52.18% 98.45% 50.11% 50.52% 98.17%
Watermark Recovery on Common Manipulations. The watermarks are represented via binary values, and the bit-wise accuracy of the recovered watermark is computed by comparing it to the original watermark following
$\textrm{ACC}(m_{\textrm{rec}},m)=\frac{l-d_{\textrm{ham}}(m_{\textrm{rec}},m)}{l},$ (10)
where $d_{\textrm{ham}}$ computes the hamming distance [43] between the recovered and original watermarks $m_{\textrm{rec}}$ and $m$, and $l$ refers to the watermark length. In the experiment, each type of manipulation in $P_{\textrm{common}}$ and $P_{\textrm{swap}}$ is performed on the watermarked image $I_{\textrm{rec}}$ before extracting the watermark via the decoder.
The watermark recovery performance in the comparative experiment is reported in Table 2. Specifically, considering the common manipulations, our method outperforms HiDDeN [61], RDA [58], CIN [28], and ARWGAN [16] at the $128\times 128$ resolution with a better average accuracy, and our performance is competitive against MBRS [18] with an acceptable average accuracy above $99\%$, achieving the second best statistic. However, it can be observed from Table 1 that MBRS heavily suffers visual quality issues with unsatisfactory PSNR and SSIM values ($33.01$ and $0.775$) in order to achieve the best bit-wise watermark recovery accuracy in Table 2. As for the $256\times 256$ resolution, we reached an average accuracy of $99.07\%$ on the common manipulations, outperforming the existing state-of-the-art methods. On the contrary, although accomplishes reasonable visual quality as listed in Table 1, MBRS experiences a large performance damping on the common manipulations in Table 2 with respect to the watermark recovery accuracy compared to the $128\times 128$ resolution version.
Watermark Recovery on Face Swapping. In this paper, the face swapping manipulation pool contains merely SimSwap [5] in the training phase, and we validated the watermark recovery performance for both in- and cross-manipulation settings by introducing other recent state-of-the-art face swapping algorithms, InfoSwap [10], UniFace [52], and E4S [27], in the testing phase. As shown in Table 2, our model consistently recovers the desired watermark messages at average accuracies of 98.02% and 96.98% for $128\times 128$ and $256\times 256$ resolutions, respectively, while HiDDeN, MBRS, CIN, and ARWGAN all obtain accuracies around 50% and RDA has accuracies below 75% against all synthetic models. Moreover, it can be concluded from the cross-manipulation tests on unseen synthetic models that by simply tuning with SimSwap, the hidden watermarks are pleasantly maintained by our method since the goals for different face swapping algorithms are identical. Meanwhile, the semi-fragile watermarking framework, FaceSigns [31], although obtains low watermark recovery accuracies as designed when facing synthetic manipulations that are expected to destroy the embedded watermarks, its performance regarding the common manipulations is unexpectedly unsatisfactory besides illustrating low visual quality as reported in Table 1, leading to an average recovery accuracy of $80.39\%$.
Discussion. In general, Salt and Pepper and Gaussian Noise are the most challenging common manipulations that cause fluctuations in the watermark recovery accuracies for all watermarking algorithms as summarized in Table 2. This is consistent with the visualization results in Fig. 2 (col. 6 & 9) such that these two manipulations add a plethora of perceivable noises to the images and thus heavily jeopardize the hidden watermarks. On the other hand, the contrastive robust watermarking frameworks fail to maintain robustness when facing the state-of-the-art face swapping models. As a result, our proposed model is the only one that maintains the bit-wise watermark recovery accuracies above 95% for all common manipulations and above 90% for all face swapping manipulations, while the contrastive methods are heavily affected by at least one of the manipulations.
Table 5: Deepfake detection performance in AUC scores against different face swapping algorithms on CelebA-HQ at $128\times 128$ and $256\times 256$ resolutions. The last row evaluates the detection performance on a mixture of all face swapping algorithms. The best performance at each resolution is in bold and the second best performance is marked with underscore.
Xception [53] SBIs [39] RECCE [4] CADDM [7] Ours
Resolution 128 256 128 256 128 256 128 256 128 256
SimSwap [5] 39.73% 71.15% 75.30% 88.94% 60.37% 69.01% 55.91% 87.66% 98.49% 98.19%
InfoSwap [10] 60.82% 65.50% 85.11% 80.50% 55.51% 52.13% 48.29% 61.39% 99.02% 99.52%
UniFace [52] 71.49% 70.34% 72.45% 79.41% 61.58% 67.35% 82.16% 82.73% 99.99% 98.99%
E4S [27] 43.40% 53.70% 63.63% 61.05% 60.88% 47.19% 64.93% 73.13% 97.71% 98.70%
Mixed 49.71% 63.04% 75.96% 80.05% 62.84% 57.13% 64.41% 72.52% 98.66% 98.97%
FF++ [38] 97.60% 90.50% 96.81% 95.57%
4.3 Cross-Dataset Experiment
We further evaluated our watermarking framework on an unseen dataset, Labeled Faces in the Wild (LFW) [14], regarding the generalization ability. Performance evaluation with respect to the visual quality is listed in Table 3. Similar results as in Table 1 can be summarized even on a different dataset. In particular, for the $128\times 128$ resolution, our approach consistently exhibits substantial performance advantages over the comparative ones. As for the $256\times 256$ resolution, MBRS [18] and our approach both have executed watermarking with promising visual qualities while FaceSigns [31] fails to do so.
Meanwhile, as listed in Table 4, the challenging manipulations concluded in Section 4.2 consistently lead to unsatisfactory bit-wise watermark recovery accuracies in this section. Consequently, although slightly affected by Salt and Pepper ($95.89\%$ and $95.99\%$) and Gaussian Blur ($99.99\%$ and $97.11\%$) at $128\times 128$ and $256\times 256$ resolutions, respectively, we successfully maintained state-of-the-art accuracies above $95\%$ for all common and face swapping manipulations.
Furthermore, obvious accuracy declines of our approach on UniFace [52] at the $256\times 256$ resolution and on E4S [27] at both resolutions can be observed in Table 2 and Table 4 on CelebA-HQ and LFW. Besides the challenges caused by the cross-manipulation setting, the former is mainly due to the setting of accepting only half-precision floating-point inputs by UniFace. Although the synthetic pipeline is unaffected, this can lead to discrepancies in the watermark recovery process at higher resolution levels because of information loss. On the other hand, the performance on E4S is affected by adopting an extra color channel, alpha, during image synthesis, which reduces the watermark feature density in the face swapped images, leading to a relatively lower watermark recovery accuracy.
4.4 Watermark Collision Check
As introduced in Section 3.2, the binary identity perceptual watermarks are derived from identity embeddings generated via well-designed face recognition tools [^2] https://github.com/serengil/deepface. To validate the claim that the identity perceptual watermarks retain the collision-free characteristic of the original identity embeddings, we performed collision checks at different watermark lengths on CelebA-HQ [19] and LFW [14]. In specific, for the sets of watermarks $M_{i}$ and $M_{j}$ of a pair of arbitrary identities $i$ and $j$, the following statement stands,
$\forall x\in M_{i}\forall y\in M_{j}(x\neq y).$ (11)
In other words, identical watermarks are not allowed to be derived from images that belong to different identities.
Table 6: Robustness ablation test on common manipulations for easier and harder parameter values that are unseen during training. The bit-wise watermark recovery accuracies are reported. Long common manipulation names are abbreviated.
Difficulty Dropout ($r$) Resize ($r$) Jpeg ($Q$) GNoise ($\mu$, $\sigma$) S&P ($r$) GBlur ($\sigma$, $k$) MBlur ($k$)
$r$ Acc $r$ Acc $Q$ Acc $\mu$, $\sigma$ Acc $r$ Acc $\sigma$, $k$ Acc $k$ Acc
Easy 0.3 99.99% 0.3 100.00% 70 100.00% 0, 0.05 100.00% 0.03 99.76% 1, 1 100.00% 1 100.00%
Regular 0.5 99.99% 0.5 100.00% 50 99.48% 0, 0.1 98.63% 0.05 95.54% 2, 3 100.00% 3 100.00%
Hard 0.7 99.46% 0.7 99.82% 30 99.38% 0, 0.2 93.24% 0.07 94.19% 3, 5 98.08% 5 99.50%
In this paper, images from CelebA-HQ with 6,217 unique identities are adopted to establish $\phi$ and produce watermarks following the pipeline in Section 3.2. Thereafter, the watermarks of images in LFW with 5,749 unseen identities are derived following the fixed rule accordingly.
As a result, there is no collision in the experiment for CelebA-HQ and LFW at both $128\times 128$ and $256\times 256$ resolutions in our collision check. Meanwhile, upon passing the collision check even though not seeing the identities in LFW when defining $\phi$, it can be concluded that the number of unique identities in CelebA-HQ is sufficient to avoid the potential bias, and the cross-dataset performance is therefore preserved.
4.5 Deepfake Detection
In this section, the Deepfake detection task is proactively achieved by evaluating the recovered watermark $m_{\textrm{rec}}$. In specific, for a target image $I$ with identity $t$ to be proactively protected with identity perceptual watermark $m$, the watermarked image $I_{\textrm{rec}}$ is firstly derived via the encoder of our framework. Then, when a face swapping algorithm from $P_{\textrm{swap}}$ is applied on $I_{\textrm{rec}}$ to generate $I_{\textrm{fake}}$ with identity $s$, the identity perceptual watermark $m_{s}$ is derived based on the image content of $I_{\textrm{fake}}$, and the robust watermark is also recovered, denoted as $m_{\textrm{rec}}$, via the decoder of our framework. In the end, the two watermarks, $m_{s}$ and $m_{\textrm{rec}}$, are compared bit-wisely and the falsification can be addressed based on their matching rate.
Since the comparative proactive watermarking approaches do not assign semantics to their watermarks, Deepfake detection cannot be accomplished without knowing the matched watermark for each image in advance, and this is usually unavailable in real-life scenarios. Meanwhile, the unsatisfactory watermark recovery accuracies in Section 4.2 suggest that the detection performance evaluation can be trivial for them. Therefore, in this experiment, we adopted several state-of-the-art passive Deepfake detectors in recent years for comparison, namely, SBIs [39], RECCE [4], and CADDM [7]. We also consider the popular milestone detector, Xception [^3] The well-trained model weights are adopted from a recent benchmark study [53].  [38], as one of the contrastive algorithms. All the models are utilized with their published weights corresponding to the optimal performance.
Since the accuracy metric requires a fixed threshold value which can lead to biases due to data distributions, we employed the area under the receiver operating characteristic (ROC) curve (AUC) score to evaluate the detection performance based on the matching rate. In a nutshell, a high AUC score demonstrates the probability that a random positive sample scores higher than a random negative sample from the testing set, in other words, the ability of the classifier to distinguish between real and fake samples. The experiment is conducted on CelebA-HQ at $128\times 128$ and $256\times 256$ resolutions by testing against each face swapping model in $P_{\textrm{swap}}$ and a mixture [^4] When conducting this experiment, samples are equally and randomly drawn for each synthetic algorithm to match up the same total quantity as the real ones. of them. To establish a balanced ratio for real and fake samples, the real images in CelebA-HQ are leveraged after randomly applying common manipulations from $P_{\textrm{common}}$.
The detection performance in AUC scores is demonstrated in Table 5. It can be observed that, although exhibit reasonable detection performance in lab-controlled scenarios, in this experiment, the passive detectors generally come up with poor AUC scores of at most 88.94%. Moreover, various statistics are observed below 70% regarding each passive model, demonstrating substantial fluctuations in the detection performance. We also listed their performance on the training dataset, FaceForensics++ [38] (FF++) dataset, for a reference. As a result, huge performance damping can be concluded for the passive detectors since the face swapping models are unseen and can generate faces with imperceptible detectable artifacts. Meanwhile, the passive detectors trained in lab-controlled scenarios fail to stay robust when facing common manipulations on the real samples that simulate the real-life scenarios.
On the contrary, our proposed approach successfully retains promising AUC scores above 97% for both seen and unseen face swapping manipulations at both resolutions. This implies that the matching rates between $m_{s}$ and $m_{\textrm{rec}}$ for images after common manipulations (real) are consistently higher than those for images after face swapping manipulations (fake). In other words, our identity perceptual watermarking framework not only robustly survives via image manipulations, but favorably distinguishes real and fake samples with respect to the changes in facial identities. As a result, for a mixture of all four manipulations including three unseen ones, our approach retains satisfactory detection performance at 98.66% and 98.97% at $128\times 128$ and $256\times 256$ resolutions, respectively, demonstrating superior generalization ability.
4.6 Ablation Study on Common Manipulations
In this paper, following the convention of steganography, we picked the common manipulations that best imitate the unavoidable image quality degradation of posting and spreading in real-life scenarios and set the parameter values with considerable difficulties. Particularly, the parameters are set equal to or harder than the early work [61, 18, 28] in our experiments. In this section, to further validate the robustness of the proposed watermarking framework, we conducted ablation testing sessions on CelebA-HQ at the $256\times 256$ resolution with easier and harder parameters that are unseen during training for the common manipulations, respectively. As a result, in addition to the outstanding performance reported in Sections 4.2 and 4.3, Table 6 implies that the robustness is completely guaranteed for common manipulations with easier parameter values, and it is also only slightly affected in harder conditions, with all accuracies still above $90\%$. In conclusion, the potential effects of common manipulations at different difficulty levels are trivial with respect to the robustness of our proposed approach.
5 Conclusion
In this paper, we propose a robust identity perceptual watermarking framework that proactively defends against malicious Deepfake face swapping. By securely assigning identity semantics to the watermarks, our approach concurrently accomplishes detection and source tracing with a single robust watermark for the first time. Meanwhile, the confidentiality is guaranteed by devising an unpredictable chaotic encryption system on the identity perceptual watermarks. Extensive experimental results have proved that our method is robust when facing both common and malicious manipulations on the watermarked images even under cross-dataset settings. Moreover, by leveraging a single Deepfake face swapping algorithm for fine-tuning, we are able to preserve promising cross-manipulation performance against unseen face swapping algorithms. The AUC scores of 98.66% and 98.97% at $128\times 128$ and $256\times 256$ resolutions further prove the promising detection ability of our approach on Deepfake face swapping, significantly outperforming the state-of-the-art passive detectors.
Admittedly, since the watermarks are identity perceptual, the proposed framework is unable to detect face re-enactment manipulations that modify facial expressions and poses but retain facial identities. Our future work will focus on designing a unified framework to proactively defend against both face swapping and face re-enactment manipulations by analyzing and introducing the corresponding facial attribute features.